OpenStudy (anonymous):

Multiply and write your answer in lowest terms: 21/26 x 39/42

OpenStudy (ikram002p):

ok so what u got till nw ?

OpenStudy (phi):

if you are good at arithmetic (memorized the multiplication table!) you can change this problem to \[ \frac{21}{26}\cdot \frac{39}{42} = \frac{3\cdot 7\cdot 3 \cdot 13}{2\cdot 13\cdot 6\cdot 7}\]

OpenStudy (phi):

notice that a 7 is in both the top and bottom, which means 7/7 which is 1 you can "cancel" the 7's from the top and bottom ask cancel the 13's you are left with \[ \frac{3 \cdot 3}{2 \cdot 6} \] notice that the 6 can be written as 2*3, so \[ \frac{3 \cdot 3}{2 \cdot 2 \cdot 3} \] we have a 3 up top and in the bottom, so cancel a pair of 3's to get \[ \frac{3}{2 \cdot 2} \]

OpenStudy (anonymous):

That was the best explanation! thank you! :)